If 3.7 moles of propane are at the temperature of 28 degrees C and are under 1.27 atm of pressure, what volume does the sample occupy.
1
Expert's answer
2020-03-26T09:28:02-0400
  1. If 3.7 moles of propane are at the temperature of 28 degrees C and are under 1.27 atm of pressure, what volume does the sample occupy.

Solution:

Given

"n = 3.7 moles"

"T=28^0C+273=300K"

"P=1.27 atm"

"R=0.0821" litre atm K-1 mol-1

"PV=nRT"

"V=nRT\/P"

"V=3.7*0.0821*300\/1.27"

"V=71.76litre"


Answer: 71.76 litre volume will be occupied